OpenStudy (clamin):

Please help.. Fill in the gaps in this arithmetic sequence: -3 , _ , _ , _ , _ , _ , 12

OpenStudy (anonymous):

-3, -0.5, 2, 4.5, 7, 9.5, 12

OpenStudy (clamin):

An arithmetic sequence has a 10th term of 17 and a 14th term of 30. Find the common difference.

OpenStudy (anonymous):

wow you said that like it was easy

OpenStudy (anonymous):

soory I don't know what a common difference is, english is not my native language

OpenStudy (clamin):

So,your answer is not right?or

OpenStudy (anonymous):

didn't you ask a second question just now?

OpenStudy (ace-n-it):

is it right

OpenStudy (clamin):

Oops my bad i thought its about the first one.

OpenStudy (anonymous):

answer for the second question should be 3.25

OpenStudy (welshfella):

10th term = a1 +9d 14th term = a1 + 13d so you need to solve the system of equation a1 + 9d = 17 a1 + 13d = 30 Subtract the first one from the second to eliminate a1 and solve for d

OpenStudy (welshfella):

@wiselunatic Please dont give direct answers Read the Code of Conduct.

OpenStudy (ace-n-it):

k